OpenStudy (anonymous):

integration ofsin^7x*cosx

OpenStudy (solomonzelman):

u= sin(x)

OpenStudy (solomonzelman):

the derivative is in there. sitting.

ganeshie8 (ganeshie8):

there is a very obvious u substitution

OpenStudy (solomonzelman):

you know though, that \(\large\color{black}{ \sin^7(x) }\) is the same thing as \(\large\color{black}{ ( \sin x)^7 }\). And that, \(\large\color{black}{ \frac{d}{dx}~\sin(x)=\cos(x) }\)

OpenStudy (solomonzelman):

you are not replying, but I'll ask you this. Have you ever done u-substitution before (because it really looks like a u-substitution problem, or are you just testing the function by "where it's coming from" saying "what could be the derivative from which this sin^7xcosx can come" ?